Cardinal Company is considering a five-year project that would require a $2,945,000 investment in equipment with a useful life of five years and no salvage value. The company’s discount rate is 18%. The project would provide net operating income in each of five years as follows:

Sales $ 2,873,000
Variable expenses 1,019,000
Contribution margin 1,854,000
Fixed expenses:
Advertising, salaries, and other fixed out-of-pocket costs $ 754,000
Depreciation 589,000
Total fixed expenses 1,343,000
Net operating income $ 511,000

__________________________________________________

2. What are the project’s annual net cash inflows?

3. What is the present value of the project’s annual net cash inflows? (Round your final answer to the nearest whole dollar amount.)

4. What is the project’s net present value? (Round discount factor(s) to 3 decimal places and final answer to the nearest whole dollar amount.)

5. What is the project profitability index for this project? (Round your answer to 2 decimal places.)
Answer #1

2) Annual Cash inflow = Net income+Depreciation = 511000+589000 = 1100000

3) Present value of annual net cash inflow = 1100000*3.127 = $3439700

4) Net present value = Present value of cash inflow-Present value of cash outflow = 3439700-2945000 = 494700

5) Profitability index = 3439700/2945000 = 1.17